OpenStudy (samsan9):

use sum or difference formulas to find the exact value of each trigonometric function. tan 23pie/12

OpenStudy (anonymous):

You want to write 23pi/12 as a sum (in this case) by finding two numbers that sum to 23 that also simplify with the 12 so that you end up with a denominator of 2, 3, 4, or 6. Can you try to do that?

OpenStudy (samsan9):

so i have to have one that has the greatest common factor?

OpenStudy (anonymous):

Start by subtracting even numbers from 23 and looking for even numbers as an answer.

OpenStudy (anonymous):

For example, 23 - 6 = 17 so that does not work because the 17 will not simplify with the 12, but 23 - 8 = 15 (whoops, not even!) and that seems to work since they both will reduce 8pi/12 + 15pi/12

OpenStudy (anonymous):

Btw, earlier I should have said "even" or a multiple of 3 since even numbers and 3 (and its multiples) will divide the 12.

OpenStudy (anonymous):

Once you reduce the 8pi/12 and the 15pi/12 you use the tangent of a sum identity. I assume you know or have it as reference.

OpenStudy (anonymous):

The identity is

OpenStudy (anonymous):

Give it a try. The answer is \[\sqrt{3}-2\]

OpenStudy (samsan9):

are you sure?

OpenStudy (anonymous):

I think so. What did you get?

OpenStudy (anonymous):

I didn't actually solve using the identity. I used http://www.wolframalpha.com/input/?i=tan%2823pi%2F12%29

OpenStudy (samsan9):

are you positive?